Find a1 if s15= 2000 and r= 3.
Answer by Boreal(15235)

s15=2000=a1*(1-3^15)/(1-3)
2000(-2)=-4000=a1(1-3^15)
a1=-4000/(1-3^15)
A1=0.00027877 to five places.
rounding to check, the 15th term is 1333, 14th 444, 13th 148, 12th 49, 11th 16 10th 5.5, 9th about 2, etc.